The conductivity of silver is 6.5 x 107per Ohm per m and number of conduction electrons per m3 is 6 x 1028. Find the mobility of conduction electrons and the drift velocity in an electric field of 1 V/m. Given m = 9.1 x 10–31 kg and e = 1.602 x 10–19 C.


1
Expert's answer
2021-06-02T09:35:50-0400

First of all, find the relaxation time:


"\tau=\frac{m_e\sigma}{ne^2}."


Now write the expression for drift velocity:


"v=\frac{eE}{m_e}·\tau=\frac{E\sigma}{ne}=6.76\text{ mm/s}."

Find the mobility of conduction electrons:


"\mu=\frac{v}{E}=6.76·10^{-3}\space\frac{\text{m}^2}{\text{V}·\text s}"